OpenStudy (anonymous):

Write the given expression as an algebraic expression in x. cos(2 cos^−1 x)

OpenStudy (anonymous):

too bad the dust is coming start with \[\cos(2\theta)=2\cos^2(\theta)-1\]

OpenStudy (anonymous):

then replace theta by \[\cos^{-1}(x)\] and you will get it in one step

OpenStudy (anonymous):

\[2x^2-1\] is your answer

OpenStudy (anonymous):

thank you!!!!!

OpenStudy (anonymous):

yw